# April 4, 2025

### v2

@@
Abstract. Televised debates between presidential candidates are often regarded as the exemplar of persuasive communication. Yet, recent evidence from Le Pennec and Pons (2023) indicates that they may not sway voters as strongly as popular belief suggests. We revisit their findings through the lens of the persuasion rate and introduce a robust framework that does not require exogenous treatment, parallel trends, or credible instruments. Instead, we leverage plausible monotonicity assumptions to partially identify the persuasion rate and related parameters. Our results reaffirm that the sharp upper bounds on the persuasive effects of TV debates remain modest.
@@ 1. Introduction
The persuasion rate, first introduced by DellaVigna and Kaplan (2007) to evaluate the efficacy of persuasive efforts, has been extensively used to make a comparison across diverse settings. Building upon this concept, Jun and Lee (2023) formulate the persuasion rate as a causal parameter that quantifies how a persuasive message affects recipients’ behavior, and they conduct a formal identification analysis based on exogenous treatment assignments or credible instruments. Extending this line of research, Jun and Lee (2024b) propose the average persuasion rate on the treated (APRT) and its reverse version (R-APRT) as additional causal parameters of interest, and they develop a difference-in-differences (DID) framework to identify, estimate, and conduct inference for them. Additional research on persuasion rates includes an analysis of persuasion types among compliers (Yu, 2025), identification of the average persuasion rate under sample selection (Possebom and Riva, 2025), an extension of the persuasion effect to continuous decisions (Kaji and Cao, 2025), and covariate-assisted bounds on the persuasion rate without monotonicity assumptions (Ji et al., 2024).
In this paper, we explore implications of two monotonicity assumptions, i.e., monotone treatment response (Manski, 1997, MTR) and monotone treatment selection (Manski and Pepper, 2000, MTS). The MTR assumption states that the effect of the informational treatment of interest is directional, and it has also been employed by Jun and Lee (2023, 2024b) in the context of the persuasion rates. In our application involving TV debates, it requires that uncertainty in vote intentions be lower after a debate takes place than before it does. Since debates provide additional information that should be helpful or at least not harmful, MTR appears plausible. For identifying the persuasion rate, MTR links the joint probability of two potential outcomes with the average treatment effect. The MTS assumption is new in studying the persuasion rate. In our context, it requires that the probability of vote choice consistency (voting exactly as intended at the time of the pre-election survey) not decrease as the survey date approaches election day, which seems natural.
@@
This robust interpretation of our estimands can be extended further. Both the persuasion rate and its reverse version are defined by conditioning on a counterfactual event. However, Pearl (1999, 2009, 2015) has advocated conditioning on an observable event rather than a counterfactual one. We show that this conceptual difference is inconsequential in our approach: our estimands serve as sharp upper bounds not only on the persuasion rate and its reverse version, but also on the probabilities of causation proposed by Pearl, when we solely rely on the MTR and MTS assumptions for identification.
@@
The rest of the paper is organized as follows. In section 2, we introduce the causal persuasion rate and related parameters, and discuss the MTR assumption, which will be used throughout. Section 3 focuses on identification, highlighting that the estimands that point-identify the parameters of interest under exogenous treatments become sharp upper bounds on those parameters when the MTS assumption is imposed. In section 4, we consider Pearl (1999)’s probabilities of causation and show that the same sharp upper bounds apply under MTR and MTS. In section 5, we apply our methodology to revisit LP’s data. In section 6, we provide concluding remarks. Finally, in the appendix, we give the proofs of our main theoretical results.
@@ 2. Preliminaries
For $d\in\{0,1\}$, let $Y(d)$ denote a binary potential outcome, $D$ the observed binary treatment, and therefore, the observed outcome is $Y=DY(1)+(1-D)Y(0)$. Because our identification results do not rely on covariates, we focus on the case without them. Alternatively, one can view our assumptions and identification results as “implicitly conditional on covariates.” The causal parameters we consider are:
@@
Assumption A imposes standard regularity. Assumption B has been used for partial identification (e.g. Manski, 1997; Manski and Pepper, 2000; Jun and Lee, 2024a) or in the context of informational treatments (Jun and Lee, 2023, 2024b). In the latter case, it is often interpreted as an assumption that says, “Persuasive messages are directional, and there is no backlash.”
In the context of our application, assumption B can be justified by LP’s research design. In their study, respondents are surveyed twice, i.e., before and after the election: the first one is to capture their voting intentions, and the second to record their final choices. Our treatment variable $D$ is equal to $1$ if the first survey has occurred after a TV debate, and it is 00 otherwise. The outcome variable $Y$, termed vote choice consistency, is set to $1$ if a respondent’s pre-election vote intention matches their eventual vote, and 00 otherwise. In particular, $Y=0$ if a respondent’s initial response is “undecided.” Formally, let $Y(1)$ and $Y(0)$ represent vote choice consistency with and without a debate, respectively. Under the MTR assumption, we have $Y(1)\geq Y(0)$, meaning a debate cannot reduce vote consistency. For the sake of intuition, suppose that all cases of vote choice inconsistency correspond to undecided voters. Then, $Y(1)=0$ and $Y(0)=1$ means that this respondent would be an undecided voter in the presence of a debate but would know exactly how to vote without the debate. This scenario is implausible if watching a debate helps voters gain information and clarify their preferences. Therefore, we exclude that possibility. A similar reasoning applies to the other instances of $Y=0$.
Under assumptions A and B, the population consists of three (nonempty) groups: (i) never-persuadable (NP): $Y(1)=0,Y(0)=0$; (ii) already-persuaded (AP): $Y(1)=1,Y(0)=1$; and (iii) treatment-persuadable (TP): $Y(1)=1,Y(0)=0$. As we never observe the pair $\{Y(1),Y(0)\}$ jointly, we cannot assign a type to each observational unit; however, we can aim to identify the share of the three groups. Using the group probabilities, we can rewrite the parameters of interest as
@@ 3. Identification
Assumption A ensures that $\mathbb{P}(Y=1\mid D=1)>0$ and $\mathbb{P}(Y=1\mid D=0)<1$, and we define the following population quantities:
@@
$\theta^{(r)}_{U}$ $=\frac{\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0)}{\mathbb{P}(Y=1\mid D=1)}.$
Both $\theta_{U}$ and $\theta^{(r)}_{U}$ are directly identified from the data. We now state the following results.
@@
Under assumptions A, C and B, we have
@@
In particular, $\theta$ and $\theta^{(r)}$ are identified by $\theta_{U}$ and $\theta^{(r)}_{U}$, respectively.
Proof.
The proof is straightforward and therefore it is omitted. ∎
Assumption C is popular, but it can be restrictive, as it precludes endogenous treatment assignment. Also, we are implicit here, but assumption C is often stated after conditioning on a sufficiently large vector of covariates, which can be too demanding in practice. To go beyond assumption C, we now consider an alternative assumption that allows endogenous treatment, namely monotone treatment selection.
@@
As an alternative, we employ the MTS assumption introduced by Manski and Pepper (2000). Formally, it requires that those in the treatment group be no less likely to “succeed” than those in the control group. Even if $D$ is a pure choice variable, assumption D stays valid as long as those who chose $D=1$ are not strictly inferior on average to those who opted out. In the context of LP, being surveyed later in the election cycle (i.e., $D=1$) naturally places respondents closer to the election day, and hence it makes vote consistency more likely, rendering MTS a plausible condition even without needing to control for covariates.
@@
Suppose that assumptions A, B and D hold. Then, the sharp bounds on $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{AP}\}$, and $\mathbb{P}\{\mathrm{TP}\}$ are given by
$\mathbb{P}\{\mathrm{NP}\}\in\left[\mathbb{P}(Y=0\mid D=1),\mathbb{P}(Y=0)\right]\quad\text{and}\quad\mathbb{P}\{\mathrm{AP}\}\in\left[\mathbb{P}(Y=1\mid D=0),\mathbb{P}(Y=1)\right].$
In particular, the sharp identified intervals on $\theta,\theta^{(r)}$ are given by
$0\leq\theta\leq\theta_{U}\quad\text{ and }\quad 0\leq\theta^{(r)}\leq\theta^{(r)}_{U}.$
@@
Theorem 1 shows that, for both $\theta$ and $\theta^{(r)}$, the same population quantities $\theta_{U}$ and $\theta^{(r)}_{U}$ arise whether we assume assumption C or assumption D, although their interpretation changes. For example, under assumptions A, C and B, $\theta$ is point-identified by $\theta_{U}$, while that same $\theta_{U}$ becomes the sharp upper bound on $\theta$ under assumptions A, B and D. Consequently, a plug-in estimator of $\theta_{U}$ is consistent for $\theta$ under exogeneity, and it will be consistent for the sharp upper bound on $\theta$ under the weaker assumption of MTS.
Also, theorem 1 establishes bounds on the shares of the NP and AP groups. In particular, $\mathbb{P}(Y=0\mid D=1)$ and $\mathbb{P}(Y=1\mid D=0)$ are now interpreted as sharp lower bounds on $\mathbb{P}\{\mathrm{NP}\}$ and $\mathbb{P}\{\mathrm{AP}\}$, respectively. From the proof of theorem 1, we can see that assumptions B and D imply
@@
where the bounds are sharp. In other words, under assumptions B and D, the mean difference (the coefficient of $D$ when regressing $Y$ on $D$) serves as a sharp upper bound on the average treatment effect, and they coincide if the treatment is exogenous. Since $\mathbb{P}\{\mathrm{TP}\}=1-\mathbb{P}\{\mathrm{NP}\}-\mathbb{P}\{\mathrm{AP}\}=\mathbb{P}\{Y(1)=1\}-\mathbb{P}\{Y(0)=1\}$, the sharp bounds on $\mathbb{P}\{\mathrm{TP}\}$ are given by
$\mathbb{P}\{\mathrm{TP}\}\in\left[0,\ \mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0)\right].$
We conclude this section with a brief discussion of estimation and inference. Plug-in estimators of $\theta_{U}$ and $\theta^{(r)}_{U}$ can be developed by regressing $Y$ on $D$ (including an intercept) and using the estimated coefficients. For inference, if assumption C is considered plausible, one can use the usual two-sided confidence intervals. Under the weaker assumption D, however, one-sided confidence intervals are more appropriate.

### v3

# February 2026

@@
Abstract.
Televised debates between presidential candidates are often regarded as the exemplar of persuasive communication. Yet, recent evidence from Le Pennec and Pons (2023) indicates that they may not sway voters as strongly as popular belief suggests. We revisit their findings through the lens of the persuasion rate and introduce a robust framework that does not require exogenous treatment, parallel trends, or credible instruments. Instead, we leverage plausible monotonicity assumptions to partially identify the persuasion rate and related parameters. Our results reaffirm that the sharp upper bounds on the persuasive effects of TV debates remain modest.
@@ 1. Introduction
The persuasion rate, first introduced by DellaVigna and Kaplan (2007) to evaluate the efficacy of persuasive efforts, has been extensively used to make a comparison across diverse settings. Building upon this concept, Jun and Lee (2023) formulate the persuasion rate as a causal parameter that quantifies how a persuasive message affects recipients’ behavior, and they conduct a formal identification analysis based on exogenous treatment assignments or credible instruments. Extending this line of research, Jun and Lee (2024b) propose the forward and backward average persuasion rates on the treated as additional causal parameters of interest, and they develop a difference-in-differences (DID) framework to identify, estimate, and conduct inference for them. Additional research on persuasion rates includes an analysis of persuasion types among compliers (Yu, 2025), identification of the average persuasion rate under sample selection (Possebom and Riva, 2025), an extension of the persuasion effect to continuous decisions (Kaji and Cao, 2025), and covariate-assisted bounds on the persuasion rate without monotonicity assumptions (Ji et al., 2024).
In this paper, we explore implications of two monotonicity assumptions: monotone treatment response (Manski, 1997, MTR) and monotone treatment selection (Manski and Pepper, 2000, MTS). The MTR assumption states that the effect of the informational treatment of interest is directional, and it has also been employed by Jun and Lee (2023, 2024b) in the context of persuasion rates. In our application to TV debates, MTR implies that exposure to the debate never increases uncertainty regarding vote intentions: i.e., debates provide information that helps voters crystallize their preferences. For identifying the persuasion rate, MTR links the joint probability of two potential outcomes with the average treatment effect. The MTS assumption is new in studying the persuasion rate. In our context, it requires that the probability of vote choice consistency (voting exactly as intended at the time of the pre-election survey) not decrease as the survey date approaches election day, which seems natural.
@@
This robust interpretation of our estimands can be extended further. Both the persuasion rate and its reverse version are defined by conditioning on a counterfactual event. However, Pearl (1999, 2009, 2015) has advocated conditioning on an observable event rather than a counterfactual one. We show that this conceptual difference is inconsequential in our approach: our estimands serve as sharp upper bounds not only on the persuasion rate and its reverse version, but also on the probabilities of causation proposed by Pearl, when we solely rely on MTR and MTS for identification.
@@
The rest of the paper is organized as follows. In Section˜2, we introduce the causal persuasion rate and related parameters, and discuss the MTR assumption, which will be used throughout. Section˜3 focuses on identification, highlighting that the estimands that point-identify the parameters of interest under exogenous treatments become sharp upper bounds on those parameters when the MTS assumption is imposed. In Section˜5, we consider Pearl (1999)’s probabilities of causation and show that the same sharp upper bounds apply under MTR and MTS. In Section˜7, we apply our methodology to revisit LP’s data. In Section˜8, we provide concluding remarks. Finally, in the appendix, we give the proofs of our main theoretical results.
@@ 2. Preliminaries
For $d\in\{0,1\}$, let $Y(d)$ denote a binary potential outcome, $D$ the observed binary treatment, and therefore, the observed outcome is Y=DY(1)+(1D)Y(0)Y=DY(1)+(1-D)Y(0). Because our identification results do not rely on covariates, we focus on the case without them. Alternatively, one can view our assumptions and identification results as “implicitly conditional on covariates.” The causal parameters we consider are:
@@
Assumption˜A imposes weak regularity conditions on the observed data. Assumption˜B has been used for partial identification (e.g. Manski, 1997; Manski and Pepper, 2000; Jun and Lee, 2024a) or in the context of informational treatments (Jun and Lee, 2023, 2024b). In the latter case, it is often interpreted as an assumption that says, “Persuasive messages are directional, and there is no backlash.”
In our application, Assumption˜B is motivated by LP’s research design, in which respondents are surveyed twice: first to elicit pre-election voting intentions, and later to record their final post-election choices. The treatment indicator $D$ equals $1$ if the initial survey occurs after a TV debate and 0 otherwise. The outcome variable $Y$, referred to as vote choice consistency, equals $1$ if a respondent’s pre-election intention matches their eventual vote, and 0 otherwise; notably, $Y=0$ whenever the initial response is “undecided.” Formally, let $Y(1)$ and $Y(0)$ represent vote choice consistency with and without a debate, respectively. Panel A of Figure˜1 illustrates the coding of these potential outcomes.
In this context, Assumption˜B means that a debate may help clarify a voter’s preference but should not confuse a previously decided voter. For instance, Scenario 1 in Panel A depicts a respondent who is undecided without the debate ($Y(0)=0$) but forms a consistent preference for Candidate A after watching it ($Y(1)=1$), a case that satisfies MTR. In contrast, Scenario 2 illustrates the reverse case, where a respondent who would have been consistent without the debate becomes undecided (and thus inconsistent) after watching it. We consider this “confusion” scenario implausible if debates serve to provide information and clarify preferences; therefore, MTR excludes this possibility. Finally, Scenarios 3 and 4 show examples where $Y(0)=Y(1)=0$, which are consistent with MTR but yield no variation in the outcome.
It is worth noting that the plausibility of the MTR assumption depends heavily on the definitions of the treatment and outcome variables. For instance, Cavgias et al. (2024) analyze the 1989 Brazilian election, where the treatment is defined as having access to a biased TV debate. In that context, the debate, which was widely regarded as biased towards the right-wing candidate, could persuade a voter to switch candidates (e.g., from Lula to Collor). Consequently, a treated voter would appear inconsistent ($Y(1)=0$) relative to their pre-debate baseline, while an untreated voter would remain consistent ($Y(0)=1$). This creates a violation of MTR. In contrast, our treatment is defined by the timing of the survey. If a debate persuades a voter to switch candidates, a respondent surveyed after the debate ($D=1$) will report their new preference, which matches their final vote ($Y(1)=1$). Meanwhile, if surveyed before the debate ($D=0$), they report their old preference, which contradicts their final vote ($Y(0)=0$). Thus, in our design, persuasion actually reinforces the MTR condition ($Y(1)\geq Y(0)$). Therefore, we view “the chance of the debate causing confusion about their own preferences” (Scenario 2) as a primary threat to our identification strategy.
Under Assumptions˜A and B, the population consists of three (nonempty) groups: (i) never-persuadable (NP): $Y(1)=0,Y(0)=0$; (ii) already-persuaded (AP): $Y(1)=1,Y(0)=1$; and (iii) treatment-persuadable (TP): $Y(1)=1,Y(0)=0$. As we never observe the pair $\{Y(1),Y(0)\}$ jointly, we cannot assign a type to each observational unit; however, we can aim to identify the share of the three groups. Using the group probabilities, we can rewrite the parameters of interest as
@@ 3. Identification
Assumption˜A ensures that $\mathbb{P}(Y=1\mid D=1)>0$ and $\mathbb{P}(Y=1\mid D=0)<1$, and we define the following population quantities:
@@
$\theta^{(r)}_{U}$ $=\frac{\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0)}{\mathbb{P}(Y=1\mid D=1)},$
which are directly identified from the data. We now state the following benchmark result; see, e.g., Jun and Lee (2023, Theorem 1).
@@
Under Assumptions˜A, C and B, we have
@@
In particular, $\theta$ and $\theta^{(r)}$ are point-identified by $\theta_{U}$ and $\theta^{(r)}_{U}$, respectively.
Assumption˜C is popular, but it can be restrictive, as it precludes endogenous treatment assignment. Also, we are implicit here, but Assumption˜C is often stated after conditioning on a sufficiently large vector of covariates, which can be too demanding in practice. To go beyond Assumption˜C, we now consider an alternative assumption that allows endogenous treatment, namely monotone treatment selection.
@@
As a robust alternative, we employ the MTS assumption introduced by Manski and Pepper (2000). Formally, it requires that those in the treatment group be no less likely to “succeed” than those in the control group. Even if $D$ is a pure choice variable, Assumption˜D stays valid as long as those who chose $D=1$ are not strictly inferior on average to those who opted out. In the context of LP, being surveyed later in the election cycle (i.e., $D=1$) naturally places respondents closer to the election day, and hence it makes vote consistency more likely, rendering MTS a plausible condition even without needing to control for covariates. See Section˜7 for more details.
@@
Suppose that Assumptions˜A, B and D hold. The sharp bounds on $\mathbb{P}\{\mathrm{NP}\}$ and $\mathbb{P}\{\mathrm{AP}\}$ are given by
$\mathbb{P}\{\mathrm{NP}\}\in\left[\mathbb{P}(Y=0\mid D=1),\mathbb{P}(Y=0)\right]\quad\text{and}\quad\mathbb{P}\{\mathrm{AP}\}\in\left[\mathbb{P}(Y=1\mid D=0),\mathbb{P}(Y=1)\right],$
where $\mathbb{P}\{\mathrm{TP}\}=1-\mathbb{P}\{\mathrm{NP}\}-\mathbb{P}\{\mathrm{AP}\}$. In particular, the sharp identified intervals on $\theta,\theta^{(r)}$ are given by
[…]
@@
Theorem˜1 shows that, for both $\theta$ and $\theta^{(r)}$, the same population quantities $\theta_{U}$ and $\theta^{(r)}_{U}$ arise whether we assume Assumption˜C or Assumption˜D, although their interpretation changes. For example, under Assumptions˜A, C and B, $\theta$ is point-identified by $\theta_{U}$, while that same $\theta_{U}$ becomes the sharp upper bound on $\theta$ under Assumptions˜A, B and D. Consequently, a plug-in estimator of $\theta_{U}$ is consistent for $\theta$ under exogeneity, and it will be consistent for the sharp upper bound on $\theta$ under the weaker assumption of MTS.
DellaVigna and Kaplan (2007) and DellaVigna and Gentzkow (2010) propose the following quantity as the persuasion rate:
$\theta_{DK}=\frac{\mathbb{P}[Y=1\mid Z=1]-\mathbb{P}[Y=1\mid Z=0]}{\mathbb{P}[D=1\mid Z=1]-\mathbb{P}[D=1\mid Z=0]}\times\frac{1}{\mathbb{P}[Y=0\mid Z=0]},$
where $Z$ is a binary instrument. Under full compliance, that is, when $D=Z$ almost surely, this quantity coincides with the upper bound on the persuasion rate derived in Theorem˜1. More generally, under partial compliance, when $D\neq Z$, it has been pointed out that $\theta_{DK}$ can be misleading and may inflate the causal persuasion rate (Jun and Lee, 2023).
In addition, Theorem˜1 establishes bounds on the shares of the NP, AP, and TP groups. In particular, $\mathbb{P}(Y=0\mid D=1)$ and $\mathbb{P}(Y=1\mid D=0)$ are now interpreted as sharp lower bounds on $\mathbb{P}\{\mathrm{NP}\}$ and $\mathbb{P}\{\mathrm{AP}\}$, respectively. Further, the proof of Theorem˜1 shows that Assumptions˜B and D yield
@@
where the bounds are sharp. The sharp bounds on the share of TP follow from here because $\mathbb{P}\{\mathrm{TP}\}=\mathbb{P}\{Y(1)=1\}-\mathbb{P}\{Y(0)=1\}$ equals the average treatment effect (ATE) under MTR.
4. Estimation and Inference
We now describe our approach to estimation and inference for the parameters of interest. Since inference for $\theta$ and $\theta^{(r)}$ follows the same principle, we will focus on the former.
4.1. Persuasion Rates
We obtain plug-in estimators for $\theta_{U}$ by using the coefficients from an OLS regression of $Y$ on $D$ with an intercept; e.g., the coefficient of $D$ corresponds to the numerator of $\theta_{U}$. If Assumption˜C is maintained, standard two-sided confidence intervals are applicable because $\theta_{U}$ point-identifies $\theta$ in this case (see Lemma˜1). Under the weaker Assumption˜D, however, $0,\theta_{U}]$ is the sharp identified bounds on $\theta$, making one-sided inference appropriate.
Specifically, under Assumption˜D, the theoretical lower bound is known to be zero. Accordingly, we construct the $(1-\alpha)$ confidence interval as
$\mathrm{CI}=\left[0,\ \hat{\theta}_{U}+c_{\alpha}\cdot\widehat{\mathrm{SE}}_{U}\right],$
where $\hat{\theta}_{U}$ is the estimator for the upper bound, $\widehat{\mathrm{SE}}_{U}$ is the corresponding standard error (computed via the delta method), and $c_{\alpha}=\Phi^{-1}(1-\alpha)$ is the standard one-sided critical value (e.g., $1.645$ for $\alpha=0.05$).
Notably, this confidence interval is valid for both the true parameter $\theta$ and the identified set $0,\theta_{U}]$. While general partial identification methods typically require larger critical values to cover the set (Imbens and Manski, 2004), our setting is simpler because the lower bound is fixed at zero. Consequently, the coverage probability for the identified set coincides with that of the parameter at the least favorable configuration (i.e., $\theta=\theta_{U}$). Thus, the standard one-sided critical value $c_{\alpha}$ ensures asymptotic coverage of at least $1-\alpha$ for both targets.
Furthermore, uniformity issues are simpler than in the general cases discussed in e.g., Canay and Shaikh (2017) too. Since $\mathbb{P}(\theta\in\mathrm{CI})\geq\mathbb{P}\{\theta_{U}(P)\in\mathrm{CI}\}$ for any data-generating process $P\in\mathcal{P}$ with the corresponding $\theta_{U}(P)\geq\theta\geq 0$, the coverage properties of $\mathrm{CI}$ will be uniform in $P\in\mathcal{P}$, provided that the normal approximation of $\{\hat{\theta}_{U}-\theta_{U}(P)\}/\widehat{\mathrm{SE}}_{U}$ is uniform. See the online appendices for more details.
4.2. Specification Testing
The identification result under Assumptions˜B and D implies that the upper bound must be non-negative ($\theta_{U}\geq 0$). This restriction is testable. Following the logic of Stoye (2009), we can construct a confidence set that becomes empty if the data provide strong evidence against the validity of the identification assumptions. Specifically, we define the modified confidence set $\mathrm{CI}_{\mathrm{mod}}$ as:
$\mathrm{CI}_{\mathrm{mod}}:=\begin{cases}[0,\ \hat{\theta}_{U}+c_{\alpha}\widehat{\mathrm{SE}}_{U}]&\text{if }\hat{\theta}_{U}+c_{\alpha}\widehat{\mathrm{SE}}_{U}\geq 0,\\ \emptyset&\text{otherwise}.\end{cases}$
The event $\mathrm{CI}_{\mathrm{mod}}=\emptyset$ corresponds to rejecting $H_{0}:\theta_{U}\geq 0$ in favor of $H_{1}:\theta_{U}<0$ at significance level $\alpha$, which indicates that MTR or MTS is violated.
4.3. Shares of Already-Persuaded and Never-Persuadable Types
As before, if Assumption˜C is maintained, standard two-sided confidence intervals are applicable because the proportions of AP and NP types are point-identified (see Lemma˜1). Under the weaker Assumption˜D, however, the situation differs from that of persuasion rates. Here, the sharp identified sets require the bounds at both ends be estimated (see Theorem˜1). To construct confidence intervals for these parameters, we adopt the method of Imbens and Manski (2004) and Stoye (2009), which adjusts the critical value to account for the estimation of the length of the identified set.
Let $\theta_{s,L},\theta_{s,U}]$ denote the identified set for type $s\in\{\mathrm{AP},\mathrm{NP}\}$. The $(1-\alpha)$ confidence interval is given by
$\mathrm{CI}_{s}=\left[\hat{\theta}_{s,L}-c_{\alpha}\widehat{\mathrm{SE}}_{s,L},\quad\hat{\theta}_{s,U}+c_{\alpha}\widehat{\mathrm{SE}}_{s,U}\right],$
where $c_{\alpha}$ is chosen to satisfy
$\Phi\left(c_{\alpha}+\frac{\hat{\Delta}_{s}}{\max(\widehat{\mathrm{SE}}_{s,L},\widehat{\mathrm{SE}}_{s,U})}\right)-\Phi(-c_{\alpha})=1-\alpha,$
with $\hat{\Delta}_{s}$ representing the estimated length of the interval. This approach ensures (uniformly) valid coverage for the true parameter $\theta_{s}$ while mitigating the conservatism inherent in intervals designed to cover the entire identified set. We refer to Stoye (2009) for formal details.